# On computational complexity and average-case hardness of shallow-depth boson sampling

### v1

@@ Abstract
Boson sampling, a computational task believed to be classically hard to simulate, is expected to hold promise for demonstrating quantum computational advantage using near-term quantum devices. However, noise in experimental implementations poses a significant challenge, potentially rendering boson sampling classically simulable and compromising its classical intractability. Numerous studies have proposed classical algorithms under various noise models that can efficiently simulate boson sampling as noise rates increase with circuit depth. To address this issue particularly related to circuit depth, we explore the viability of achieving quantum computational advantage through boson sampling with shallow-depth linear optical circuits. Specifically, as the average-case hardness of estimating output probabilities of boson sampling is a crucial ingredient in demonstrating its classical intractability, we make progress on establishing the average-case hardness confined to logarithmic-depth regimes. We also obtain the average-case hardness for logarithmic-depth Fock-state boson sampling subject to lossy environments and for the logarithmic-depth Gaussian boson sampling. By providing complexity-theoretical backgrounds for the classical simulation hardness of logarithmic-depth boson sampling, we expect that our findings will mark a crucial step towards a more noise-tolerant demonstration of quantum advantage with shallow-depth boson sampling.
preprint: APS/123-QED
@@ I Introduction
Boson sampling is a computational task that is complexity-theoretically proven to be hard to classically simulate under plausible assumptions [1, 2, 3]. Accordingly, boson sampling has gathered significant attention, as it would possibly play a key role in the experimental demonstration of quantum computational advantage using near-term quantum devices. However, the implementation of boson sampling in experimental settings with near-term quantum devices is inevitably subject to various sources of noise [4, 5, 6, 7]. The problem is that those noises would possibly rule out the classical intractability of boson sampling, and thus potentially hinder the experimental demonstration of quantum advantage with boson sampling. Indeed, both for finite-size near-term experiments and asymptotic limits as system size scales, numerous studies [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] have proposed efficient classical simulation algorithms of boson sampling under various noise models, such as photon loss, partial distinguishability, gaussian noise, etc. Their results indicate that as the noise rate of boson sampler increases, it eventually renders such a noisy sampler classically simulable. Moreover, as the noise is typically accumulated with each circuit depth, the quantum signal for classical intractability exhibits exponential decay with increasing circuit depth. Hence, circuits with polynomially increasing depth with system size would suffer from significantly enlarged noise rates, posing substantial challenges to achieving quantum advantage in such settings.
@@ I.1 Outlines: average-case hardness of shallow-depth boson sampling
We set our goal as proving the hardness of classical simulation of boson sampling in the shallow-depth regime, specifically for approximate simulation within total variation distance error. Two key ingredients for the current hardness proof of the approximate simulation of boson sampling are $(i)$ average-case #P-hardness of output probability approximation up to sufficiently large additive imprecision $\epsilon$, and $(ii)$ hiding property. Informally, average-case hardness means that approximating the output probability of boson sampling with high probability over randomly chosen circuits (i.e., on average over circuits) is #P-hard. Here, by choosing random circuit instances that have the hiding property (i.e., symmetric over outcomes), one can reduce the average-case instances for the hardness from circuit instances to outcome instances, which is a crucial step to prove the hardness of approximate simulation within total variation distance error (See Appendix A for more details).
@@
Most of the current theoretical foundations of the average-case hardness of boson sampling rely on global Haar random unitary circuits [1, 2, 3, 20, 21, 22], as they almost satisfy the two conditions described above. Namely, the outcome instances can be effectively hidden by global Haar random unitaries, and approximating the output probability within sufficiently large $\epsilon$ on average over global Haar random circuit instances is #P-hard under some conjectures. However, the problem is that implementing global Haar random unitary requires at least polynomial circuit depth (e.g., see [23, 24]), and thus not implementable in sub-polynomial circuit depths. Accordingly, the hardness results built upon global Haar random unitaries cannot be directly applied to the shallow-depth boson sampling we are interested in, necessitating a different approach from them.
[…]
@@
To deal with those problems we take the following approach: first, we consider shallow linear optical circuit architectures composed of geometrically non-local gates. In fact, implementation of non-local gates is promising for near-term experimental settings; for example, experiments of linear optical systems based on trapped ions [30, 31] and photonic architecture [6] implemented long-range interactions. Also, since we cannot implement global Haar random unitary within shallow-depth regime, we instead employ local random circuit ensemble for random circuit instances in shallow circuit architecture, inspired by the hardness results of random circuit sampling [32, 33, 21, 34, 35]. Here, local random distribution in this context means that each gate composing the circuit is independently chosen Haar random gate; we note that recent experimental setups of boson sampling [4, 5, 6, 7] also follow such circuit distribution, but with geometrically local architectures.
However, local random distribution poses a subsequent challenge toward the average-case hardness of shallow-depth boson sampling, that is, the absence of the hiding property. Since the output symmetry of boson sampling over local random circuit distribution is not evident, the random outcome instances cannot be efficiently hidden by random circuit instances. This means that even if we find the average-case hardness over randomly chosen circuits for a fixed outcome, it still does not lead to the classical simulation hardness grounded in Stockmeyer’s reduction from the average-case approximation over randomly chosen outcomes [1]. To address this issue, we prove the average-case hardness over both the randomly chosen outcome and randomly chosen circuit, by establishing a worst-to-average-case reduction for both outcome and circuit instances. Specifically, our reduction process is composed of two steps: $(i)$ from a given fixed outcome to a randomly chosen collision-free outcome, and $(ii)$ from a worst-case circuit to a randomly chosen circuit over local random circuit distribution.
@@
To sum up, we show the average-case hardness over outcomes and circuit instances for shallow circuit architectures composed of non-local gates and employing the local random circuit ensemble. We informally present here our average-case hardness result of boson sampling in the logarithmic depth regime, for photon number $N$ and mode number $M\propto N^{\gamma}$ with $\gamma\geq 1$.
@@
There exists a $O(\log N)$-depth linear optical circuit architecture such that approximating output probability of boson sampling within additive error $2^{-O(N^{\gamma+1}(\log N)^{2})}$ with high probability over randomly chosen circuits in the circuit architecture for high probability over randomly chosen collision-free outcomes is #$\rm{P}$-hard under $\rm{BPP}^{\rm{NP}}$ reduction.
Also, since our average-case hardness result considers both the random outcomes and random circuits due to the absence of the hiding property, it is not straightforward to show the classical simulation hardness of shallow-depth boson sampling as in the original boson sampling proposal [1]. Accordingly, we show how our average-case hardness result over both the randomly chosen outcomes and circuits leads to the classical simulation hardness argument. This implies that improving the additive imprecision for our average-case hardness result is the only remaining problem for the fully theoretically guaranteed classical intractability of shallow-depth boson sampling.
@@
If the allowed additive imprecision for the problem in Theorem 1 to be #$\rm{P}$-hard is improved to $\epsilon=2^{-(\gamma-1)N\log N-O(N)}$, the approximate boson sampling for the shallow-depth circuit in Theorem 1, up to constant total variation distance, is classically hard to simulate.
Now we provide an outline of our results, which is depicted in Fig. 1. We first define in Sec. II a shallow-depth circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ composed of non-local gates, which we will use throughout our results. Next, in Sec. III we prove the worst-case #P-hardness of approximating output probability $p_{\bm{s}}(C)$ of a fixed outcome $\bm{s}$ of boson sampling, for any circuit $C$ in the shallow circuit architecture previously defined. In Sec. IV we prove the average-case #P-hardness of approximating output probability $p_{\bm{s}}(U)$ for randomly chosen outcome $\bm{s}$ and randomly chosen circuit $U$ in the shallow circuit architecture, by establishing worst-to-average-case reduction. We prove in Sec. V how our average-case hardness results over both the random outcomes and random circuits lead to the classical simulation hardness. We also extend our average-case hardness result to the Gaussian boson sampling scheme in Sec. VI, and to the lossy boson sampling subject to photon loss channels in Sec. VII. In Sec. VIII, we conclude with several remarks.
@@
II Notations
Let us define the total mode number as $M$, where we set $M$ as a power of 2 for simplicity. We set the output photon number $N$ polynomially related to $M$ as $M=c_{0}N^{\gamma}$, for a constant $c_{0}$ and $\gamma\geq 1$ satisfying $M\geq 2N$. We use the notation $\bm{s}$ as an $M$-dimensional output configuration vector for the collision-free outcome, such that each element $s_{i}$ of $\bm{s}$ denotes photon number in $i$th mode. Namely, $\bm{s}=(s_{1},\dots,s_{M})$ where each $s_{i}\in\{0,1\}$ with $\sum_{i=1}^{M}s_{i}=N$, so that the number of possible configurations of $\bm{s}$ is $\binom{M}{N}$. We define $p_{\bm{s}}(C)$ as an output probability of a linear optical circuit (unitary) matrix $C$ for the outcome $\bm{s}$ from a predefined input configuration $\bm{t}$. For collision-free input and output, $p_{\bm{s}}(C)$ can be represented as [1]
$p_{\bm{s}}(C)=|\text{Per}(C_{\bm{s},\bm{t}})|^{2}$ (1)
where $C_{\bm{s},\bm{t}}$ is a $N$ by $N$ matrix obtained by taking $s_{i}$ copies of the $i$th row and $t_{j}$ copies of the $j$th column of the matrix $C$.
We note that an $M$-mode linear optical circuit can be represented by an $M$ by $M$ unitary matrix in U$(M)$ which unitarily transforms $M$ mode operators. Specifically, we can represent a single-mode gate (i.e., a phase shifter) as a U$(1)$ matrix to the mode, and a two-mode gate (i.e., a beam splitter) as a U$(2)$ matrix along the modes. Also, the parallel application of gates can be represented as a unitary matrix with a block matrix form, and the serial application of gates can be represented as matrix multiplication of the unitary matrices. Accordingly, throughout this work, we will interchangeably use the terminology ‘(linear optical) circuit’ and ‘(unitary) matrix’.
We first define the linear optical circuit architecture, for a more rigorous analysis of the hardness proof.
@@ II Notations
For the shallow-depth circuit architecture, specifically in logarithmic depth, we define the shallow linear optical circuit architecture of circuit depth $D=\log M$, using the convention used in [36, 37].
Definition 2.
We illustrate in Fig. 2 the circuit architecture $\mathcal{B}$ and $\mathcal{B}^{*}$, and the form of their corresponding unitary matrix. Next, we define the Kaleidoscope circuit architecture proposed in [36], using the butterfly circuit architecture defined above.
Definition 3.
We define Kaleidoscope circuit architecture $\mathcal{B}\mathcal{B}^{*}$ as a serial application of $\mathcal{B}$ over $\mathcal{B}^{*}$. We also define $q$-Kaleidoscope circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ as a repeat of the Kaleidoscope circuit architecture, with repetition number $q\in\mathbb{N}$.
Here, the circuit depth of $q$-Kaleidoscope architecture is $D=2q\log M$, which is indeed a logarithmic depth in $N$ for $q=O(1)$. Throughout this paper, we will focus on the $q$-Kaleidoscope circuit architecture with $q=O(1)$ to demonstrate the hardness results for shallow-depth circuits. One motivation for employing this linear optical circuit architecture is that it enjoys a useful property that is crucial for our analysis; Ref. [36] shows that for $M$ a power of $2$, any $M$ mode permutation circuit can be implemented within $\mathcal{B}\mathcal{B}^{*}$.
@@
Let $\bm{P}$ be an arbitrary $M\times M$ permutation matrix with $M$ a power of 2. Then $\bm{P}$ can be efficiently implemented in $\mathcal{B}\mathcal{B}^{*}$ using two-mode permutation gates, i.e., $\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ and $\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$.

### v2

@@ Abstract
Boson Sampling, a computational task believed to be classically hard to simulate, is expected to hold promise for demonstrating quantum computational advantage using near-term quantum devices. However, noise in experimental implementations poses a significant challenge, potentially rendering Boson Sampling classically simulable and compromising its classical intractability. Numerous studies have proposed classical algorithms that can efficiently simulate Boson Sampling under various noise models, particularly as noise rates increase with circuit depth. To address this challenge, we investigate the viability of achieving quantum computational advantage through Boson Sampling implemented with shallow-depth linear optical circuits. In particular, as the average-case hardness of estimating output probabilities of Boson Sampling is a crucial ingredient in demonstrating its classical intractability, we make progress on establishing the average-case hardness of Boson Sampling confined to logarithmic-depth regimes. We also obtain the average-case hardness for logarithmic-depth Fock-state Boson Sampling subject to lossy environments and for the logarithmic-depth Gaussian Boson Sampling. By providing complexity-theoretical backgrounds for the classical simulation hardness of logarithmic-depth Boson Sampling, we expect that our findings will mark a crucial step towards a more noise-tolerant demonstration of quantum advantage with shallow-depth Boson Sampling.
@@ I Introduction
Boson Sampling is a computational task that is complexity-theoretically proven to be hard to classically simulate under plausible assumptions [1, 2, 3]. Accordingly, Boson Sampling has gathered significant attention, as it would possibly play a key role in the experimental demonstration of quantum computational advantage using near-term quantum devices. However, the implementation of Boson Sampling in experimental settings with near-term quantum devices is inevitably subject to various sources of noise [4, 5, 6, 7]. The problem is that those noises would possibly rule out the classical intractability of Boson Sampling, and thus potentially hinder the experimental demonstration of quantum advantage with Boson Sampling. Indeed, both for finite-size near-term experiments and asymptotic limits as system size scales, numerous studies [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26] have proposed efficient classical simulation algorithms of Boson Sampling under various noise models, such as photon loss, partial distinguishability, random Gaussian noise, etc. Their results indicate that as the noise rate of boson sampler increases, it eventually renders such a noisy sampler classically simulable. Moreover, as the noise is typically accumulated with each circuit depth, the quantum signal for classical intractability exhibits exponential decay with increasing circuit depth. Hence, circuits with polynomially increasing depth with system size would suffer from significantly enlarged noise rates, posing substantial challenges to achieving quantum advantage in such settings.
@@ I.1 Outlines: average-case hardness of shallow-depth Boson Sampling
We set our goal as proving the hardness of classical simulation of Boson Sampling in the shallow-depth regime, specifically for approximate simulation within total variation distance error. Two key ingredients for the current hardness proof of the approximate simulation of Boson Sampling are $(i)$ average-case #P-hardness of output probability approximation up to sufficiently large additive imprecision $\epsilon$, and $(ii)$ hiding property. Informally, average-case hardness means that approximating the output probability of Boson Sampling with high probability over randomly chosen circuits (i.e., on average over circuits) is #P-hard. Here, by choosing random circuit ensembles that have the hiding property (i.e., symmetry over outcomes), one can reduce the average-case instances for the hardness from circuit instances to outcome instances, which is a crucial step to prove the hardness of approximate simulation within total variation distance error (See Appendix A for more details).
@@
Most of the current theoretical foundations of the average-case hardness of Boson Sampling rely on global Haar random unitary circuits [1, 2, 3, 27, 28, 29, 30], as they almost satisfy the two conditions described above. Namely, the outcome instances can be effectively hidden by global Haar random unitaries, and approximating the output probability within sufficiently large $\epsilon$ on average over global Haar random circuit instances is #P-hard under some conjectures. However, the problem for our case is that implementing a global Haar random unitary requires at least polynomially large circuit depth (e.g., see [31, 32]), and thus is not implementable in sub-polynomial circuit depths. This poses a challenge in proving the average-case hardness of Boson Sampling in the shallow-depth regime—specifically, the lack of the hiding property. Due to the absence of hiding in the shallow-depth regime, the random outcome instances of Boson Sampling cannot be hidden by random circuit instances. Accordingly, even if we establish average-case hardness over randomly chosen circuits for a fixed outcome, it does not directly imply classical simulation hardness via Stockmeyer’s reduction, which requires average-case hardness over randomly chosen outcomes [1].
Another problem is that there already exist efficient classical algorithms that can approximately simulate shallow-depth Boson Sampling in certain circumstances, which directly rule out the classical simulation hardness in the shallow-depth regime for those cases. While exact simulation of Boson Sampling is classically hard even for constant-depth circuits [33], approximate simulation becomes easy for 1-dimensional local log-depth circuits [34, 35] and also for more general dimension local circuits under some constraints [36, 37]. According to their results, if we use circuits composed of only geometrically local gates, a polynomial circuit depth is generally required for a sufficiently large correlation to achieve approximate simulation hardness. Those results indicate that we cannot expect the hardness results in the most general case of shallow-depth circuits composed of local gates only.
@@
To deal with those problems, we take the following approach: First, we consider shallow linear optical circuit architectures that fully employ geometrically non-local gates. In fact, the implementation of non-local gates is promising for near-term experimental settings; for example, experiments of linear optical systems based on trapped ions [38, 39] and photonic architecture [6] implemented long-range interactions. Because of the absence of hiding, we derive the average-case hardness over both randomly chosen outcomes and randomly chosen circuits over the shallow circuit architectures, by establishing a worst-to-average-case reduction for both outcome and circuit instances. In short, this reduction process can be done in two steps: $(i)$ from a given fixed outcome to a randomly chosen collision-free outcome, and $(ii)$ from a worst-case circuit to a randomly chosen circuit over a local random circuit distribution.
For random circuit ensembles, we employ a local random circuit ensemble in the shallow circuit architecture, inspired by the hardness results of random circuit sampling [40, 41, 28, 42, 43]. Here, the local random distribution in this context means that each gate composing the circuit is independently chosen Haar random gate; we note that recent experimental setups of Boson Sampling [4, 5, 6, 7] follow a similar circuit distribution, but with geometrically local architectures. Additionally, due to the absence of symmetry, the local random circuit ensemble itself does not guarantee the bosonic birthday paradox, which is crucial for the hardness of Boson Sampling when using only collision-free outcomes [1]. Hence, we also include a random permutation circuit at the input of our random circuit to guarantee the bosonic birthday paradox for our random circuit ensemble (see Appendix B). In fact, this random permutation circuit can be implemented within the shallow depth regime (see, e.g., Lemma 1), and can be experimentally realized by preparing a random input configuration instead.
@@
To sum up, we show the average-case hardness over outcomes and circuit instances for shallow circuit architectures composed of geometrically non-local gates and employing the local random circuit ensemble. We informally present here our average-case hardness result of Boson Sampling in the logarithmic depth regime, for photon number $N$ and mode number $M\propto N^{\gamma}$ with $\gamma\geq 2$.
@@
There exists an $O(\log N)$-depth linear optical circuit architecture $\mathcal{A}$ composed of $O(N^{\gamma}\log N)$ number of geometrically non-local gates, such that approximating output probability of Boson Sampling within additive error $2^{-O(N^{\gamma+1}(\log N)^{2})}$ with high probability over the outcomes and circuits in $\mathcal{A}$ is #$\rm{P}$-hard under $\rm{BPP}^{\rm{NP}}$ reduction.
Also, since our average-case hardness result regards both the random outcomes and random circuits, it is not straightforward to show the classical simulation hardness of shallow-depth Boson Sampling as in the original Boson Sampling proposal [1]. Accordingly, we show how our average-case hardness result over both the randomly chosen outcomes and circuits leads to the classical simulation hardness argument. This implies that improving the additive imprecision for our average-case hardness result is the only remaining problem for the fully theoretically guaranteed classical intractability of shallow-depth Boson Sampling.
@@
For $M\propto N^{\gamma}$ with $\gamma\geq 2$, if the allowed additive imprecision for the problem in Theorem 1 to be #$\rm{P}$-hard is improved to $\epsilon=2^{-(\gamma-1)N\log N-O(N)}$, the approximate Boson Sampling for the shallow-depth circuit in Theorem 1, up to constant total variation distance, is classically hard to simulate.
Now we provide an outline of our results, which is depicted in Fig. 1. We first define in Sec. II a shallow-depth circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ composed of non-local gates, which we will use throughout our results. Next, in Sec. III, we prove the worst-case #P-hardness of approximating output probability $p_{\bm{s}}(C)$ of a fixed outcome $\bm{s}$ of Boson Sampling, for any circuit $C$ in the shallow circuit architecture previously defined. In Sec. IV, we prove the average-case #P-hardness of approximating output probability $p_{\bm{s}}(U)$ for randomly chosen outcome $\bm{s}$ and randomly chosen circuit $U$ in the shallow circuit architecture, by establishing worst-to-average-case reduction. We prove in Sec. V how our average-case hardness results over both the random outcomes and random circuits lead to the classical simulation hardness. We also extend our average-case hardness result to the Gaussian Boson Sampling scheme in Sec. VI, and to the lossy Boson Sampling subject to photon loss channels in Sec. VII. In Sec. VIII, we conclude with several remarks.
@@
II Our setting
II.1 Notations
Let us define the total mode number as $M$, where we set $M$ as a power of 2 for simplicity. We set the output photon number $N$ polynomially related to $M$ as $M=c_{0}N^{\gamma}$ for a constant $c_{0}$ and $\gamma\geq 2$, to ensure that collision-free outcomes dominate the probability weight over all possible $N$-photon outcomes over $M$ modes (i.e., by bosonic birthday paradox; see Refs. [1, 44] and Appendix B). We use the notation $\bm{s}$ as an $M$-dimensional output configuration vector representing $N$-photon outcome over $M$ modes, such that each element $s_{i}$ ($i\in[M]$) of $\bm{s}$ denotes the occupied number of photons in the $i$th mode. Namely, $\bm{s}$ can be expressed as $\bm{s}=(s_{1},\dots,s_{M})$ where each $s_{i}\in[N]$ with the constraint $\sum_{i=1}^{M}s_{i}=N$. Also, let $\bm{t}$ be an $M$-dimensional input configuration vector corresponding to an $N$-photon input state over $M$ modes, such that $\bm{t}=(t_{1},\dots,t_{M})$ where now each $t_{i}\in\{0,1\}$ with $\sum_{i=1}^{M}t_{i}=N$.
Hereafter, we use the terminology Boson Sampling representing Boson Sampling with Fock-state inputs as in the original Boson Sampling proposal [1]; we will explicitly use the terminology Gaussian Boson Sampling for Boson Sampling with Gaussian state inputs [2]. We now define $p_{\bm{s}}(C)$ as the output probability of Boson Sampling with a linear optical circuit (unitary) matrix $C$ to obtain an outcome $\bm{s}$ from a fixed input configuration $\bm{t}$. More precisely, $p_{\bm{s}}(C)$ can be represented as [1]
$p_{\bm{s}}(C)=\frac{1}{\prod_{i=1}^{M}s_{i}!}|\text{Per}(C_{\bm{s},\bm{t}})|^{2},$ (1)
where $C_{\bm{s},\bm{t}}$ is an $N$ by $N$ matrix obtained by taking $s_{i}$ copies of the $i$th row and $t_{j}$ copies of the $j$th column of the matrix $C$. Here, we omit the $\bm{t}$ dependence in the output probability $p_{\bm{s}}(C)$ for simplicity because we will fix the input configuration throughout the paper.
We note that an $M$-mode linear optical circuit can be represented by an $M$ by $M$ unitary matrix in U$(M)$, which unitarily transforms mode operators. Specifically, we can represent a single-mode gate (i.e., a phase shifter) as a U$(1)$ matrix to the mode, and a two-mode gate (i.e., a beam splitter) as a U$(2)$ matrix along the modes. Also, the parallel application of gates can be represented as a unitary matrix with a block matrix form, and the serial application of gates can be represented as matrix multiplication of the unitary matrices. Accordingly to this correspondence, throughout this work, we will interchangeably use the terminology ‘(linear optical) circuit’ and ‘(unitary) matrix’.
We first define the linear optical circuit architecture for a more rigorous analysis of the hardness proof.
@@ II.1 Notations
Given the definition of the circuit architecture $\mathcal{A}$, we denote $m$ as the number of gates in the circuit architecture. Also, we denote $D$ as the depth of the circuit architecture, where we define a unit depth as a single step of parallel application of gates, such that any serial application of gates is not allowed in the unit depth. Here, we emphasize that single-mode gates represented by U(1) matrices do not contribute to the depth count since they can be absorbed by the nearest two-mode gates represented by U(2) matrices.
In this work, we consider the shallow-depth circuit architecture, specifically in logarithmic depth regime $D=O(\log N)$. As we only consider the binary logarithm throughout this work, we will omit the base of the logarithm such that $\log(\cdot)$ implies $\log_{2}(\cdot)$ hereafter. We now define the shallow linear optical circuit architecture of circuit depth $D=\log M$ with the number of gates $m=\frac{M}{2}\log M$, using the convention used in [45, 46].
Definition 2 (Butterfly circuit architecture).
@@
We emphasize that in Definition 2, gates are applied in parallel for all $j=1,2,\dots,2^{D-L}$ and $k=1,2,\dots,2^{L-1}$. Hence, by definition of the circuit depth, the total depth for $\mathcal{B}^{*}$ (and thus also $\mathcal{B}^{*}$) is given by $D=\log M$. We illustrate in Fig. 2 the circuit architecture $\mathcal{B}$ and $\mathcal{B}^{*}$ for $162^{4}=$ 16 modes (such that the circuit depth is $D=4$ each), and the form of their corresponding unitary matrix.
Next, we define the Kaleidoscope circuit architecture proposed in [45], using the butterfly circuit architecture defined above.
Definition 3 (Kaleidoscope circuit architecture).
We define Kaleidoscope circuit architecture $\mathcal{B}\mathcal{B}^{*}$ as a serial application of $\mathcal{B}$ over $\mathcal{B}^{*}$ in Definition 2. We define $q$-Kaleidoscope circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ as a repeat of the Kaleidoscope circuit architecture, for a repetition number $q\in\mathbb{N}$.
Here, the circuit depth of the $q$-Kaleidoscope architecture is $D=2q\log M$, which is indeed a logarithmic depth $D=O(\log N)$ for $q=O(1)$. Also, the number of gates for the $q$-Kaleidoscope architecture is $m=qM\log M=O(N^{\gamma}\log N)$ for $q=O(1)$. Note that, because all the gates are geometrically non-local except for the gates in the first and the last depth for each $\mathcal{B}\mathcal{B}^{*}$, the number of non-local gates also scales $O(N^{\gamma}\log N)$ for $q=O(1)$. Throughout this work, we will focus on the $q$-Kaleidoscope circuit architecture with $q=O(1)$ to demonstrate the hardness results for shallow-depth circuits. The motivation for employing this linear optical circuit architecture is that it enjoys a useful property that is crucial for our analysis, which is that for $M$ a power of $2$, any $M$-mode permutation circuit can be implemented within $\mathcal{B}\mathcal{B}^{*}$:
@@
Let $\bm{P}$ be an arbitrary $M\times M$ permutation matrix with $M$ a power of 2. Then $\bm{P}$ can be efficiently implemented in $\mathcal{B}\mathcal{B}^{*}$ using $M\log M$ number of two-mode permutation gates, i.e., $\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ and $\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$.
II.2 Definitions of random circuit ensemble and random outcome ensemble
In this work, we show the average-case hardness over both randomly chosen outcomes and randomly chosen circuits by establishing a worst-to-average-case reduction for both outcomes and circuit instances. Here, for the average-case instances, it is crucial to choose a proper circuit ensemble and outcome ensemble to successfully establish the worst-to-average-case reduction.
For the random circuit instances, we consider the product of two random circuit ensembles. The first is the local random circuit ensemble, which is the circuit ensemble with each gate being independently drawn from the local Haar measure.
Definition 4 (Local random circuit ensemble).
Indeed, this circuit ensemble is similar to the random circuit ensemble employed by the previous random circuit sampling proposals [40, 41, 28, 42, 43] which employ gate-wise Haar random unitary circuits. Next, we introduce our second random circuit ensemble, which is the random permutation circuit ensemble over the fixed shallow-depth circuit architecture $\mathcal{B}\mathcal{B}^{*}$.
Definition 5 (Random permutation circuit ensemble).
We define $\mathcal{P}$ as the uniform distribution over $M$-mode permutation circuits over the fixed Kaleidoscope circuit architecture $\mathcal{B}\mathcal{B}^{*}$ composed of $M\log M$ number of two-mode permutation gates $\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ and $\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$. Each $\bm{P}\sim\mathcal{P}$ represents an $M$ by $M$ permutation matrix, uniformly chosen from $M!$ number of possible permutation matrices.
Note that this random permutation circuit ensemble can be implemented in the shallow-depth circuit architecture $\mathcal{B}\mathcal{B}^{*}$ by Lemma 1. Before proceeding, we briefly remark on the motivation for including the permutation circuit ensemble in our random circuit construction. Specifically, for any $M$-mode circuit architecture $\mathcal{A}$, including $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ considered in this work, a circuit drawn from the product circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$ guarantees bosonic birthday paradox. The bosonic birthday paradox ensures an upper bound on the probability of obtaining collision outcomes over all possible outcomes of Boson Sampling (see Appendix B for more details). More precisely, for a randomly chosen circuit $U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}$, the total probability weight of collision outcomes in the Boson Sampling output distribution is upper bounded by $O(\frac{N^{2}}{M})$. Indeed, this bound becomes sufficiently small in the dilute regime $M=\Omega(N^{2})$, which is the primary regime of interest in this work. Accordingly, the bosonic birthday paradox ensures that collision-free outcomes dominate the probability weight for our setup, and thus enables us to confine the outcome space only to collision-free outcomes. Also, it is worth emphasizing that, from an experimental perspective, the setup for our random circuit ensemble can be realized by constructing the local random circuit $\mathcal{H}_{\mathcal{A}}$ and preparing a random input configuration over the experiments.
For the random outcome instances, as we can now only focus on collision-free outcomes for the hardness argument, we also define the random outcome ensemble as the uniform distribution over collision-free outcomes.
Definition 6 (Random collision-free outcome ensemble).
We define $\,\mathcal{G}_{M,N}$ as the uniform distribution over $\binom{M}{N}$ number of collision-free outcomes of Boson Sampling with $M$ modes and $N$ photons. Each outcome $\bm{s}\sim\mathcal{G}_{M,N}$ is an $M$-dimensional output configuration vector representing the collision-free outcome, such that $\bm{s}=(s_{1},\dots,s_{M})$ where each $s_{i}\in\{0,1\}$ with $\sum_{i=1}^{M}s_{i}=N$.